# A Survey of Model Architectures in Information Retrieval

### v2

@@ 7 Large Language Models for IR
The natural evolution from pre-trained encoders is the recent ascendance of Large Language Models (LLMs). [^4] The term “Large Language Model” lacks a precise, universally accepted definition in the literature. In this survey, we use the term to refer to models with over one billion parameters that are pre-trained with a text generation objective, such as text infilling (e.g., T5) or causal language modeling (e.g., the GPT series), and are often optionally post-trained for instruction following and human preference alignment. While building on the same transformer principles discussed in §˜6, the sheer scale and generative capabilities of modern instruction-following LLMs are reshaping the architectural landscape of IR. These models are not just larger backbones for feature extraction; their proficiency in language understanding, generation, and instruction-following allows them to take on entirely new roles. Trained to align with human preferences (OpenAI, 2023; Gemini et al., 2023; Bai et al., 2022), LLMs can perform complex tasks such as reasoning (Wei et al., 2022; Hurst et al., 2024; Guo et al., 2025), tool usage (Schick et al., 2023; Patil et al., 2024b; Qin et al., 2024a; Patil et al., 2024a) and planning (Song et al., 2023; Huang et al., 2024a). In this section, we review how these powerful models — spanning decoder-only and encoder-decoder architectures — are being adapted for IR tasks, moving beyond established paradigms into new frontiers of retrieval, reranking, and direct generation.
LLM as Retriever.
@@
A straightforward yet highly effective application of LLMs is to serve as the backbone for bi-encoder retrieval models. The dramatic increase in parameter count and training data provides LLMs with richer world knowledge and a more nuanced understanding of semantics compared to their smaller BERT-sized predecessors. This directly translates to performance improvements. Neelakantan et al. (2022) fine-tuned a series of off-the-shelf GPT models towards text and code representation. They empirically verified that the bi-encoder retriever’s performance can benefit from increased backbone language model capacity. Muennighoff (2022); Ma et al. (2024c) empirically verified the effectiveness of scaling the size of dense retrievers with open-weight models such as GPT-J (Wang & Komatsuzaki, 2021) and Llama-2 (Touvron et al., 2023). Today, common text retrieval benchmarks like BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2023) are dominated by proprietary and open-source LLM-based embedding models (Wang et al., 2023a; Li et al., 2023d; Lee et al., 2025; Zhang et al., 2025; Muennighoff et al., 2025, inter alia).
@@
A parallel line of research focuses on adapting the unidirectional architecture of decoder-only LLMs (e.g., Llama) to better suit the needs of bidirectional text representation. Standard decoder-only models are optimized for next-token prediction, which may not be ideal for creating a single summary vector for a whole text. To address this, LLM2Vec (BehnamGhader et al., 2024) enables bidirectional attention and further trains Llama-2 (Touvron et al., 2023) with specific unsupervised and supervised adaptive tasks. Similarly, NV-Embed (Lee et al., 2025) introduces a new latent attention mechanism to produce improved representations, leading to improved performance on the MTEB benchmark compared to directly enabling bi-directional attention. GritLM (Muennighoff et al., 2024) finetunes Mistral family models with both dense retrieval task and text generation task with different attention mechanisms and demonstrate the potential of unifying retrieval and generation with one single foundation model. We show a shortlist of works that use LLM as retriever backbones in Table˜4.
LLM as Reranker.
@@
The reranking task has also been significantly advanced by LLMs, which introduce new capabilities beyond the cross-encoder architecture discussed in §˜6. This evolution can be categorized into two main architectural approaches: fine-tuning and zero-shot prompting. We show a shortlist of works that use LLM as rerankers in Table˜5.
@@
First, LLMs can be fine-tuned as powerful rerankers. Extending earlier work on BERT-type models, researchers have applied similar techniques to larger encoder-decoder models like T5 (Raffel et al., 2020) and decoder-only models like Llama (Touvron et al., 2023). Nogueira et al. (2020) fine-tuned T5 with a classification loss, treating reranking as a binary relevance decision. RankT5 (Zhuang et al., 2023a) took a more direct approach by fine-tuning T5 to output a numerical relevance score, optimizing the model with established ranking losses like RankNet (Burges, 2010). Further, Zhuang et al. (2023a) also investigated the impact of language model architectures (T5 encoder-decoder versus T5 encoder), loss functions (pointwise, pairwise, listwise), and pooling strategies to ranking performance. Instead of scoring documents individually, ListT5 (Yoon et al., 2024) adopts a Fusion-in-Decoder architecture (Izacard & Grave, 2021) to process and rank an entire list of candidate documents in a single forward pass. More specifically, the architecture consists of an encoder and decoder, where the encoder takes a query and multiple passages as input in parallel, and the decoder outputs a sorted list of input passages in the decreasing order of relevance, achieving better reranking efficiency compared to pointwise methods.
@@
The long-context capabilities of modern LLMs have also enabled a new reranking paradigm. RankLlama (Ma et al., 2024c) demonstrated superior pointwise reranking performance compared to BERT and T5-based rerankers for long document reranking where the input is truncated at 4,096 tokens. Yang et al. (2025b) finetuned QwQ-32B Team (2025) for listwise reranking where multiple passages are concatenated together as input and achieved better performance than pointwise reranking, leveraging the reasoning and long context capability of the strong base model.
@@
Second, the instruction-following ability of modern LLMs has unlocked zero-shot reranking via prompting. This paradigm requires no task-specific fine-tuning. Instead, the LLM is prompted with a query and a list of candidate documents and asked to identify the most relevant ones (Ma et al., 2023; Zhang et al., 2023b; Pradeep et al., 2023b; c; Sun et al., 2023). This listwise approach is a natural fit for the long context windows of models like GPT-4. To mitigate the high computational cost and context length limitations of processing full documents, Liu et al. (2024b) proposed using passage embeddings as compact document representations for the LLM, training a specialized reranker that operates on these embeddings to improve efficiency. As this line of research primarily involves prompt engineering rather than architectural changes, we refer readers to a recent survey (Zhu et al., 2023) for further details.
Generative Retrieval.
@@
Perhaps the most radical architectural shift enabled by LLMs is generative retrieval. Traditional IR systems follow the "index-retrieval-rank" paradigm (Schütze et al., 2008; Xu et al., 2025b). Generative retrieval fundamentally challenges this by reframing retrieval as a sequence-to-sequence task. Instead of searching an index, an autoregressive language model is trained to directly generate the unique identifiers (DocIDs) of relevant documents in response to a query.
The foundational work in generative retrieval emerged from the entity linking domain with Generation of ENtity REtrieval (GENRE, De Cao et al., 2021). Rather than treating entity retrieval as a classification problem over atomic labels with dense representations, GENRE reframed it as a generative task where an encoder-decoder model produces entity names autoregressively, token-by-token, conditioned on the input context. This autoregressive formulation addressed several key limitations of traditional entity retrieval approaches. First, it enabled direct cross-encoding between context and entity names, capturing fine-grained interactions that simple vector dot products could miss. Second, the memory footprint was greatly reduced because model parameters scaled with vocabulary size rather than entity count, eliminating the need to store dense representations for large entity sets. Third, the exact softmax loss could be computed efficiently without requiring negative sampling during training. During inference, GENRE used constrained beam search to generate entity names sequentially. While GENRE could theoretically operate without candidate retrieval, it achieved optimal performance when high-quality candidate lists were available to constrain the generation space (Lai et al., 2022).
Building on GENRE’s success in entity linking, DSI introduced generative retrieval to the broader document retrieval domain (Tay et al., 2022). The core innovation was fully parameterizing traditional "index-retrieve" pipelines within a single neural model, where all corpus information is encoded in the model parameters rather than external indices (Tay et al., 2022; Pradeep et al., 2023a). DSI operates through two fundamental sequence-to-sequence tasks that can be trained jointly or sequentially (He et al., 2024). The indexing task (Learn to Index) trains the model to associate document content with corresponding document identifiers (Kishore et al., 2023). The retrieval task (Learn to Retrieve) trains the model to generate relevant document identifiers in response to user queries. A critical design choice in DSI involves document identifier representation. Early work explored various approaches including atomic identifiers (unique integers), simple string identifiers (titles, URLs), and semantically structured identifiers created through clustering algorithms (Chen et al., 2023a). The semantically structured approach, which applies hierarchical clustering to group documents by their semantic embeddings, proved most effective (Zhu et al., 2023). During inference, DSI employs constrained beam search over a prefix tree (trie) constructed from all document identifiers to ensure only valid identifiers are generated. This constraint mechanism prevents the model from generating non-existent document identifiers while maintaining the autoregressive generation process.
DSI can be implemented in two variants: classification-based approaches that use a classification layer to output atomic document identifiers, and generative approaches that autoregressively generate identifier strings (Mehta et al., 2022). The T5 model backbone (Raffel et al., 2020) serves as the foundation for most DSI implementations, trained with cross-entropy loss on both indexing and retrieval objectives. However, DSI faces significant challenges in dynamic environments. The tight coupling between index and retrieval modules makes updating the corpus computationally expensive, typically requiring full model retraining (Mehta et al., 2022). The technical implementation of generative retrieval systems centers on sequence-to-sequence modeling, where models are trained using cross-entropy loss to learn mappings between queries/documents and their corresponding identifiers (Tay et al., 2022; Ren et al., 2023). The model autoregressively generates identifier tokens with conditional probability, where each token depends on the query and previously generated tokens.
A critical technical requirement is ensuring only valid document identifiers are generated during inference. This is typically achieved through constrained beam search over a prefix tree (trie) constructed from all valid document identifiers (Tang et al., 2023). Alternative approaches include constrained greedy search and FM-index structures (Tang et al., 2023). Document identifier design represents a fundamental implementation challenge. Early approaches explored atomic identifiers (unique integers), simple string identifiers (titles, URLs), and semantically structured identifiers created through clustering algorithms (Chen et al., 2023b). More sophisticated methods have introduced n-gram-based identifiers (Bevilacqua et al., 2022) and hierarchical representations using techniques like residual quantization (Zeng et al., 2024). Scalability poses a major challenge for generative retrieval systems. Most research has focused on relatively small collections with several hundred thousand documents, as the approach struggles to scale to very large corpora. The memory and computational requirements grow substantially as corpus size increases. Generative retrieval is an active and rapidly evolving research area; we direct interested readers to a dedicated survey (Li et al., 2025c) for a comprehensive review.
@@
RepLlama (Ma et al., 2024c) LLM Encoder Llama Listwise Loss
@@
LLaRA (Li et al., 2023a) LLM Encoder Llama Adaptation + Contrastive Training
@@
Qwen3-Retriever (Zhang et al., 2025) LLM Encoder Qwen3 Synthetic Data + Listwise Loss
@@
monoT5 (Nogueira et al., 2020) LM T5 Classification
Nogueira dos Santos et al. (2020) LM BART Unlikelihood
QLM-T5 (Zhuang et al., 2021) LM T5 Language Modeling
duoT5 (Pradeep et al., 2021) LM T5 Pairwise Loss
RankT5 (Zhuang et al., 2023a) LLM Encoder + Prediction Layer T5 Listwise Loss
@@
SGPT-CE (Muennighoff, 2022) LM GPT-J & GPT-Neo Listwise Loss
RankLlama (Ma et al., 2024c) LLM Encoder + Prediction Layer Llama Listwise Loss
RankMamba (Xu, 2024) LLM Encoder + Prediction Layer Mamba Listwise Loss
RankVicuna (Pradeep et al., 2023b) LM Vicuna Listwise
RankZephyr (Pradeep et al., 2023c) LM Zephyr Listwise
Zhang et al. (2023b) LM Code-LLaMA-Instruct Listwise
Liu et al. (2024b) Embedding + LM Mistral Listwise
Qwen3-Reranker (Zhang et al., 2025) LM Qwen3 Synthetic Data + Pairwise Loss
@@
Zhuang et al. (2023b) LM Multiple Pointwise Prompting
Zhuang et al. (2024a) LM Flan-PaLM-S Pointwise Prompting
UPR (Sachan et al., 2022) LM T5 & GPT-Neo Pointwise Prompting
PRP (Qin et al., 2024b) LM Flan-UL2 Pairwise Prompting
Yan et al. (2024) LM Flan-UL2 Pairwise Prompting
Zhuang et al. (2024b) LM Flan-T5 Pairwise & Setwise Prompting
LRL (Ma et al., 2023) LM GPT-3 Listwise Prompting
RankGPT-3.5 (Sun et al., 2023) LM GPT-3.5 Listwise Prompting
RankGPT-4 (Sun et al., 2023) LM GPT-4 Listwise Prompting
Table 5: Summary of IR model architecture utilizing large language models for reranking. Nogueira dos Santos et al. (2020) and Zhuang et al. (2021) revisit the statistic language model problem with modern transformer-based models, including BART (Lewis et al., 2020a) T5 (Raffel et al., 2020) and GPT-2 (Radford et al., 2019).
Broader Ecosystem and Concluding Remarks
Beyond core architectural changes, LLMs are influencing the entire IR ecosystem. Their advanced generative and understanding capabilities are being harnessed for crucial supporting tasks. For instance, modern IR systems require extensive labeled data for training, which is expensive to create. A promising line of work uses LLMs to synthesize high-quality training data (queries, relevant passages, and hard negatives), a technique known as "in-pars" generation and its variants (Bonifacio et al., 2022; Boytsov et al., 2024; Dai et al., 2023; Lee et al., 2024; Mo et al., 2024a; c). From an evaluation perspective, LLMs’ language understanding has led to research on using them as proxies for human relevance judges, which could dramatically accelerate the evaluation cycle (Faggioli et al., 2023; 2024; Clarke & Dietz, 2024). We also point readers to a comprehensive survey on conversational information retrieval (Mo et al., 2024b), another area being reshaped by LLMs.

### v3

@@ 7 Large Language Models for IR
The natural evolution from pre-trained encoders is the recent ascendance of Large Language Models (LLMs). [^4] The term “Large Language Model” lacks a precise, universally accepted definition in the literature. In this survey, we use the term to refer to models with over one billion parameters that are pre-trained with a text generation objective, such as text infilling (e.g., T5) or causal language modeling (e.g., the GPT series), and are often optionally post-trained for instruction following and human preference alignment. While building on the same transformer principles discussed in Section˜6, the sheer scale and generative capabilities of modern instruction-following LLMs are reshaping the architectural landscape of IR. These models are not just larger backbones for feature extraction; their proficiency in language understanding, generation, and instruction-following allows them to take on entirely new roles. Trained to align with human preferences (OpenAI et al., 2023; Gemini Team Google et al., 2023; Bai et al., 2022), LLMs can perform complex tasks such as reasoning (Wei et al., 2022; OpenAI et al., 2024; Guo et al., 2025), tool usage (Schick et al., 2023; Patil et al., 2024b; Qin et al., 2024a; Patil et al., 2024a) and planning (Song et al., 2023; Huang et al., 2024a). In this section, we review how these powerful models — spanning decoder-only and encoder-decoder architectures — are being adapted for IR tasks, moving beyond established paradigms into new frontiers of retrieval, reranking, and direct generation.
7.1 LLM as Retriever
A straightforward yet highly effective application of LLMs is to serve as the backbone for bi-encoder retrieval models. We categorize these developments into backbone scaling, architectural adaptation, and unified modeling. We show a shortlist of works in Table˜5.
Scaling Bi-Encoders.
@@ 7.1 LLM as Retriever
The dramatic increase in parameter count and training data provides LLMs with richer world knowledge and a more nuanced understanding of semantics compared to their smaller BERT-sized predecessors. This directly translates to performance improvements. Neelakantan et al. (2022) fine-tuned a series of off-the-shelf GPT models towards text and code representation. They empirically verified that the bi-encoder retriever’s performance can benefit from increased backbone language model capacity. Muennighoff (2022); Ma et al. (2024b) empirically verified the effectiveness of scaling the size of dense retrievers with open-weight models such as GPT-J (Wang & Komatsuzaki, 2021) and Llama-2 (Touvron et al., 2023). Today, common text retrieval benchmarks like BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2023) are dominated by proprietary and open-source LLM-based embedding models (Wang et al., 2023a; Li et al., 2023d; Lee et al., 2025; Zhang et al., 2025b; Muennighoff et al., 2025, inter alia).
Decoder-Only Adaptation.
@@
A parallel line of research focuses on adapting the unidirectional architecture of decoder-only LLMs (e.g., Llama) to better suit the needs of bidirectional text representation. Standard decoder-only models are optimized for next-token prediction, which may not be ideal for creating a single summary vector for a whole text. To address this, LLM2Vec (BehnamGhader et al., 2024) enables bidirectional attention and further trains Llama-2 (Touvron et al., 2023) with specific unsupervised and supervised adaptive tasks. Similarly, NV-Embed (Lee et al., 2025) introduces a new latent attention mechanism to produce improved representations, leading to improved performance on the MTEB benchmark compared to directly enabling bi-directional attention.
Unified Modeling.
GritLM (Muennighoff et al., 2025) fine-tunes Mistral family models with both dense retrieval task and text generation task with different attention mechanisms and demonstrate the potential of unifying retrieval and generation with one single foundation model.
7.2 LLM as Reranker
@@
The reranking task has also been significantly advanced by LLMs, which introduce new capabilities beyond the cross-encoder architecture discussed in Section˜6. This evolution can be categorized into two main architectural approaches: fine-tuning and zero-shot prompting. We show a shortlist of works that use LLM as rerankers in Table˜6.
Fine-Tuned Rerankers.
First, LLMs can be fine-tuned as powerful rerankers. Extending earlier work on BERT-type models, researchers have applied similar techniques to larger encoder-decoder models like T5 (Raffel et al., 2020) and decoder-only models like Llama (Touvron et al., 2023).
@@
  • Pointwise and Pairwise: Nogueira et al. (2020) fine-tuned T5 with a classification loss, treating reranking as a binary relevance decision. RankT5 (Zhuang et al., 2023a) took a more direct approach by fine-tuning T5 to output a numerical relevance score, optimizing the model with established ranking losses like RankNet (Burges, 2010). Further, Zhuang et al. (2023a) also investigated the impact of language model architectures (T5 encoder-decoder versus T5 encoder), loss functions (pointwise, pairwise, listwise), and pooling strategies to ranking performance.
  • Listwise: Instead of scoring documents individually, ListT5 (Yoon et al., 2024) adopts a Fusion-in-Decoder architecture (Izacard & Grave, 2021) to process and rank an entire list of candidate documents in a single forward pass. More specifically, the architecture consists of an encoder and decoder, where the encoder takes a query and multiple passages as input in parallel, and the decoder outputs a sorted list of input passages in the decreasing order of relevance, achieving better reranking efficiency compared to pointwise methods. Yang et al. (2025b) fine-tuned QwQ-32B (Qwen Team, 2025) for listwise reranking where multiple passages are concatenated together as input and achieved better performance than pointwise reranking, leveraging the reasoning and long context capability of the strong base model.
@@
  • Long-Context Reranking: The long-context capabilities of modern LLMs have also enabled a new reranking paradigm. RankLlama (Ma et al., 2024b) demonstrated superior pointwise reranking performance compared to BERT and T5-based rerankers for long document reranking where the input is truncated at 4,096 tokens.
Zero-Shot / Few-Shot Prompting.
@@
Second, the instruction-following ability of modern LLMs has unlocked zero-shot and few-shot reranking via prompting. This paradigm requires no task-specific fine-tuning. Instead, the LLM is prompted with a query and a list of candidate documents and asked to identify the most relevant ones (Ma et al., 2023; Zhang et al., 2023c; Pradeep et al., 2023b; c; Sun et al., 2023). This listwise approach is a natural fit for the long context windows of models like GPT-4. To mitigate the high computational cost and context length limitations of processing full documents, Liu et al. (2024b) proposed using passage embeddings as compact document representations for the LLM, training a specialized reranker that operates on these embeddings to improve efficiency. As this line of research primarily involves prompt engineering rather than architectural changes, we refer readers to a recent survey (Zhu et al., 2023) for further details.
7.3 Generative Retrieval
@@
Perhaps the most radical architectural shift enabled by LLMs is generative retrieval. Traditional IR systems follow the “retrieve-then-rerank” paradigm (Schütze et al., 2008; Asai et al., 2024a; Xu et al., 2025c). Generative retrieval fundamentally challenges this by reframing retrieval as a sequence-to-sequence task. Instead of searching an index, an autoregressive language model is trained to directly generate the unique document identifiers (DocIDs) of relevant documents in response to a query.
Evolution of the Paradigm.
The foundational work in generative retrieval emerged from the entity linking domain with Generation of ENtity REtrieval (GENRE, De Cao et al., 2021). Rather than treating entity retrieval as a classification problem over atomic labels with dense representations, GENRE reframed it as a generative task where an encoder-decoder model produces entity names autoregressively, token-by-token, conditioned on the input context. Building on GENRE’s success, DSI introduced generative retrieval to the broader document retrieval domain (Tay et al., 2022). The core innovation was fully parameterizing traditional “retrieve-then-rerank” pipelines within a single neural model, where all corpus information is encoded in the model parameters rather than external indices (Tay et al., 2022; Pradeep et al., 2023a). DSI operates through two fundamental sequence-to-sequence tasks that can be trained jointly or sequentially (He et al., 2024): the indexing task (Learn to Index) and the retrieval task (Learn to Retrieve).
Table 4: Taxonomy of identifier types in Generative Retrieval. The choice of identifier is a key architectural distinction.
Identifier Type Description
Atomic Identifiers Unique integers assigned to documents. Simple but lacks semantic generalization.
String Identifiers Natural language strings (titles, URLs). Leverages pre-trained knowledge but can be ambiguous.
Semantic Identifiers Structured IDs derived from clustering embeddings. Enables semantic generalization in the ID space.
Taxonomy of Identifiers.
A critical design choice in DSI involves document identifier representation (summarized in Table˜4).
  • Atomic and String Identifiers: Early work explored atomic identifiers (unique integers) and simple string identifiers (titles, URLs) (Chen et al., 2023a). DSI can be implemented in two variants: classification-based approaches that use a classification layer to output atomic document identifiers, and generative approaches that autoregressively generate identifier strings (Mehta et al., 2023). More sophisticated methods have introduced n-gram-based identifiers (Bevilacqua et al., 2022).
  • Semantic Identifiers: Semantically structured identifiers created through clustering algorithms proved most effective (Zhu et al., 2023). This often involves hierarchical representations using techniques like residual quantization (Zeng et al., 2024), where the model learns to associate document content with corresponding semantically meaningful document identifiers (Kishore et al., 2023).
Inference and Constraints.
The technical implementation of generative retrieval systems centers on sequence-to-sequence modeling. A critical technical requirement is ensuring only valid document identifiers are generated during inference. This is typically achieved through constrained beam search over a prefix tree (trie) constructed from all valid document identifiers (Tang et al., 2023). Alternative approaches include constrained greedy search and FM-index structures (Tang et al., 2023). The T5 model backbone (Raffel et al., 2020) serves as the foundation for most DSI implementations, trained with cross-entropy loss on both indexing and retrieval objectives.
Challenges.
Generative retrieval faces significant challenges in dynamic environments. The tight coupling between index and retrieval modules makes updating the corpus computationally expensive, typically requiring full model retraining (Mehta et al., 2023). Scalability poses a major challenge; most research has focused on relatively small collections, as the memory and computational requirements grow substantially as corpus size increases. Generative retrieval is an active and rapidly evolving research area; we direct interested readers to a dedicated survey (Li et al., 2025d) for a comprehensive review.
@@
RepLlama (Ma et al., 2024b) LLM Encoder Llama Listwise Loss
@@
LLaRA (Li et al., 2023b) LLM Encoder Llama Adaptation + Contrastive Training
@@
Qwen3-Retriever (Zhang et al., 2025b) LLM Encoder Qwen3 Synthetic Data + Listwise Loss
Table 6: Summary of IR model architecture utilizing large language models for reranking. Nogueira dos Santos et al. (2020) and Zhuang et al. (2021) revisit the statistic language model problem with modern transformer-based models, including BART (Lewis et al., 2020a), T5 (Raffel et al., 2020), and GPT-2 (Radford et al., 2019). We use Seq2Seq LLM to refer to encoder-decoder architecture language models such as T5 and BART, and Casual LLM to refer to modern LLMs with causal language model architecture like GPT family models.
@@
monoT5 (Nogueira et al., 2020) Seq2Seq LM T5 Classification
Nogueira dos Santos et al. (2020) Seq2Seq LLM BART Unlikelihood
QLM-T5 (Zhuang et al., 2021) Seq2Seq LLM T5 Language Modeling
duoT5 (Pradeep et al., 2021) Seq2Seq LLM T5 Pairwise Loss
RankT5 (Zhuang et al., 2023a) Seq2Seq LLM Encoder + Prediction Layer T5 Listwise Loss
@@
SGPT-CE (Muennighoff, 2022) Causal LLM GPT-J & GPT-Neo Listwise Loss
RankLlama (Ma et al., 2024b) Casual LLM Encoder + Prediction Layer Llama Listwise Loss
RankMamba (Xu, 2024) Causal LLM Encoder + Prediction Layer Mamba Listwise Loss
RankVicuna (Pradeep et al., 2023b) Causal LLM Vicuna Listwise
RankZephyr (Pradeep et al., 2023c) Causal LLM Zephyr Listwise
Zhang et al. (2023c) Causal LLM Code-LLaMA-Instruct Listwise
Liu et al. (2024b) Embedding + Causal LLM Mistral Listwise
Qwen3-Reranker (Zhang et al., 2025b) Causal LLM Qwen3 Synthetic Data + Pairwise Loss
@@
Zhuang et al. (2023b) Causal LLM Multiple Pointwise Prompting
Zhuang et al. (2024a) Causal LLM Flan-PaLM-S Pointwise Prompting
UPR (Sachan et al., 2022) Seq2Seq LM & Causal LLM T5 & GPT-Neo Pointwise Prompting
PRP (Qin et al., 2024b) Seq2Seq LM Flan-UL2 Pairwise Prompting
Yan et al. (2024) Seq2Seq LM Flan-UL2 Pairwise Prompting
Zhuang et al. (2024b) Seq2Seq LM Flan-T5 Pairwise & Setwise Prompting
LRL (Ma et al., 2023) Casual LLM GPT-3 Listwise Prompting
RankGPT-3.5 (Sun et al., 2023) Causal LLM GPT-3.5 Listwise Prompting
RankGPT-4 (Sun et al., 2023) Causal LLM GPT-4 Listwise Prompting
7.4 Broader Ecosystem and Concluding Remarks
Beyond core architectural changes, LLMs are influencing the entire IR ecosystem. Their advanced generative and understanding capabilities are being harnessed for crucial supporting tasks:
  • Data Synthesis: Modern IR systems require extensive labeled data for training, which is expensive to create. A promising line of work is to use LLMs to synthesize high-quality training data (e.g., queries, relevant passages, and hard negatives) (Bonifacio et al., 2022; Boytsov et al., 2024; Dai et al., 2023; Lee et al., 2024; Mo et al., 2024a; c; Zhang et al., 2025b).
  • Evaluation: From an evaluation perspective, LLMs’ language understanding has led to research on using them as proxies for human relevance judges, which could dramatically accelerate the evaluation cycle (Faggioli et al., 2023; 2024; Clarke & Dietz, 2024).
We also point readers to a comprehensive survey on conversational information retrieval (Mo et al., 2024b), another area being reshaped by LLMs.